OpenStudy (anonymous):

Which set of shapes contains members that are always similar to one another? trapezoids isosceles triangles equilateral triangles rectangles pentagons @lovepurple @StudyGurl14

OpenStudy (anonymous):

oh hi

OpenStudy (anonymous):

lol hi

OpenStudy (anonymous):

i went to your profile and gave you some medals from you answering questions

OpenStudy (anonymous):

lol i know i saw

OpenStudy (anonymous):

lol i told you i would ll keep my word

OpenStudy (studygurl14):

Equilateral

OpenStudy (anonymous):

okay

OpenStudy (studygurl14):

All equilateral triangles have three 60 degree angles, no matter what the side length, so all equilateral triangles are similar to each other

OpenStudy (anonymous):

equalitiral augh what big wordy

OpenStudy (anonymous):

lololol

OpenStudy (anonymous):

ABC has side lengths of 10 units, 20 units, and 24 units. XYZ is similar to ABC and its longest side is 60 units in length. What is the perimeter of XYZ? 162 units 150 units 135 units 125 units 90 units

OpenStudy (anonymous):

yeah for real i puased everytime i typed a letter

OpenStudy (anonymous):

lol i do that somtimes

OpenStudy (anonymous):

lolollol

OpenStudy (anonymous):

i think 90 or 125

OpenStudy (studygurl14):

\(\large\frac{60}{24}=\frac{x}{20}\rightarrow1200=24x\rightarrow x=50\) \(\large\frac{60}{24}=\frac{y}{10}\rightarrow600=24x\rightarrow y=25\) 25 + 50 + 60 = 135

OpenStudy (anonymous):

i was thinking 125

OpenStudy (studygurl14):

It's 135

OpenStudy (anonymous):

ooh well 135

OpenStudy (anonymous):

lol okay

OpenStudy (anonymous):

The figure shows eight congruent triangles made by dividing a square that has an area of 64 cm2. What is the area of ABH? 20 cm2 16 cm2 8 cm2 6 cm2 4 cm2

OpenStudy (anonymous):

how many more questions @AfromanNiggA

OpenStudy (anonymous):

just 1 more

OpenStudy (anonymous):

this one to be exact

OpenStudy (anonymous):

k

OpenStudy (studygurl14):

8 i think

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

8

OpenStudy (anonymous):

medal